# ProGeny II: the impact of libraries and model configurations on inferred galaxy properties in SED fitting

### v1

@@ 4.1 Impact on derived stellar masses
The relative deviation in derived stellar masses caused by the choice of SPL is indicated in Fig. 4, where for each SPL we present the mean offset in stellar mass relative to ProGeny. FSPS produces the largest stellar masses, followed closely by ProGeny and CB19, whose masses are only $\sim$0.05 dex lower on average. The other SPLs tend to produce notably lower stellar masses, offset by $\sim$0.1-0.15 dex on average. This behaviour can be directly linked to the underlying stellar recycling fraction implemented in each SPL. As demonstrated by Robotham & Bellstedt (in prep), the recycled fraction as a function of stellar age varies between the SPLs. A larger fraction of the remaining stellar mass is accounted for in SPLs with more stellar remnants. FSPS, CB19, and ProGeny all have similar remaining fractions with stellar age, hence explaining why these SPLs tend to recover the most stellar mass. Jones et al. (2022) compared properties of galaxies derived using CB16 (the early form of CB19) and BPASS, and found the CB16 stellar masses to be greater than those from BPASS by 0.14 dex. This is consistent with the offset that we derive between CB19 and BPASS. The open points in Fig. 4 compare the total stellar mass formed throughout the history of the galaxy, which does not account for mass loss due to stellar evolution. It can be seen that the total mass formed by BPASS is more consistent with the other SPLs than the remaining mass, highlighting that the origin of the mass offset comes from the recycled fraction in BPASS.
Figure 5: The median stellar mass offsets derived when using different SED fitting configurations, as compared with our fiducial configuration. The shaded region indicates the deviation of the masses caused by the choice of SPL, where a wider grey bar indicates that the SPL produces more discrepant properties.
@@
The shaded region in Fig. 5 shows the typical mass deviation caused when selecting each SPL, relative to others. What is notable, is that the mass offsets caused by almost all SED configuration differences is less than the offsets simply caused by the choice of SPL.
@@
Figure 6: Stellar mass offsets for individual galaxies using different SED configurations, all with the ProGeny SPL. Top panel shows the impact of assuming a constant metallicity, the middle panel shows the impact of fixing the metallciity to solar, and the bottom panel shows the impact of assuming a stepwise parametric SFH.
While Fig. 5 shows the median stellar mass impact, it is important to note that these discrepancies are stellar mass dependent. While the overall offset created by the constant Z assumption is always less than 0.01 dex, at stellar masses $M_{*}>10^{10.5}\rm{M}_{\odot}$ the values are overestimated by 0.06 dex in BPASS, and underestimated by just over 0.04 dex in both CB19 and ProGeny. Subpanels (p)–(r) of Fig. 6 shows the offsets of individual galaxies using the ProGeny SPL for each SED configuration, as compared with the fiducial implementation. The impact of ignoring the presence of a relation between a galaxy’s stellar mass and metallicity is evident in sub-panel (q), where the impact of fixing the metallicity to solar is a very strong mass dependence of the stellar mass offset. At low stellar masses, galaxies are forcibly modelled with metallicity values that are too large, forcing the model to produce younger stars to create a sufficiently blue SED. This results in a lower mass-to-light, and therefore the masses are underestimated. At high stellar masses however the assumption of a solar metallicity is more realistic, and hence the stellar mass offset is minimal. Sub-panel (r) of Fig. 6 also shows that while the overall effect of a stepwise versus functional parametrisation is to overestimate the stellar masses by around 0.1 dex in ProGeny, at the highest masses the values are actually consistent. These trends are qualitatively similar for all SPLs except BPASS, where the impact of a constant metallicity is reversed at higher stellar masses (see sub-panel (l) of Fig. 6).
@@ 4.2 Impact on derived SFR
Presenting the results from just the fiducial configuration, we compare the impact of selecting different SPLs on the derived star formation rates (SFRs) relative to the ProGeny-derived SFR in Fig. 7. As with Fig. 4, the shaded region shows a range of 0.1 dex, and each black point shows the median offset between two SPLs. Values are generally consistent within ±$\pm 0.1$ dex, where M05 produces the highest values, and BPASS produces the lowest.
Figure 8: The median SFR offsets derived when using different SED fitting configurations, as compared with our fiducial configuration. The shaded region indicates the deviation of the values caused by the choice of SPL. All SFR measurements less than $10^{-6}\,\rm{M}_{\odot}\rm{yr}^{-1}$ are shifted to $10^{-6}\,\rm{M}_{\odot}\rm{yr}^{-1}$ for the sake of this comparison (as the SFR measurements derived using massfunc_snorm_trunc can be arbitrarily small).
The relative impact of different ProSpect implementation configurations as compared with the impact of SPL selection is then shown in Fig. 8. As in Fig. 5, the coloured points represent the offset caused by the SED fitting configuration relative to the fiducial configuration, whereas the dark shaded region represents the mean absolute deviation (MAD) of the SPL values for the fiducial run. For runs where the coloured point lies outside of the shaded region, this is an indication that the SED configuration has a greater impact on the SFR measurements than the choice of SPL. The SPL choice is the dominant impact on SFRs in almost all cases. The only exceptions to this are the $Z_{\odot}$ runs conducted with BPASS and M05, where the resulting SFRs are measured to be around 0.2 dex higher than the fiducial implementation, which is greater than the typical impact from SPL choice.
@@
A parameter space that has been extremely well-studied in the literature is the stellar mass – SFR plane, which presents the co-called Star Formation Main Sequence (SFMS). Fig. 9 shows how the fiducial implementation produces varying SFMS derivations using the six SPLs. The recovery of the SFMS is generally consistent, with a very similar normalisation, and a similar bending of the MS at high masses. This is reflected by the coloured lines, which represent the running median of all galaxies that have SFRs greater than 1 dex below the SFMS fit by Thorne et al. (2021). While these lines are not meant to be a recovery of the SFMS, they provide a straightforward way of comparing the overall behaviour in the bottom panel. There is a population of high-mass “quenching” galaxies that sit just below the MS in all outputs except that of CB19. These galaxies consistently occupy this region, suggesting that the constraint of this lower SFR is real. There is also a population of low-mass “bursting” galaxies produced by M05, however the fact that they are not recovered by any other SPL suggests that the separation to higher SFRs is an artefact of M05. The other SPLs do recover these sources consistently as highly star-forming galaxies, in that they either sit directly on or above the main sequence defined by Thorne et al. (2021). Finally, we note that the derived SFMS all consistently show greater scatter in the b5 SFH implementation than the fiducial implementation, highlighting the consequence of the freedom in the present-day star formation rate when using stepwise bins in the SFH implementation.
@@ 4.3 Impact on the derived metallicity
The derived mass–metallicity relations using different SSPs is given in Fig. 10. Each panel represents the output derived using a different SPL, where values for individual galaxies are indicated as grey points, with the running median shown in the coloured line. The inherent limit in metallicity range provided by each SPL can be seen in this figure, where ProGeny has the widest range in possible metallicity values, whereas M05 has the most truncated range. For a comparison with observations of direct metallicity measurements, we show the MZR from Tremonti et al. (2004) as the dashed black line in each panel. The bottom right subpanel compares the median MZR derived by all SPLs. The overall shape for each SPL is similar, recovering the classic increase of metallicity with stellar mass, where there is a break in the relation at $\sim 10^{10}\rm{M}_{\odot}$, and a flattening of the curve at the highest stellar masses.
It is notable that the low-mass values for metallicity are recovered quite differently for the various SPLs. BC03, FSPS, CB19 and ProGeny all have lower median values of metallicity of $\sim 3\times 10^{-3}$. Metallicity values for low-mass galaxies in M05 and BPASS are substantially lower, with $<10^{-3}$. For M05 this is largely driven by the fact that low-mass galaxies are hitting the lower bound of the available templates, while BPASS seems to genuinely prefer fitting lower metallicity values. The SPL capable of producing a MZR most similar in shape to that of Tremonti et al. (2004) is BC03.
@@
Figure 12: The mean Zfinal offsets derived when using different SED fitting configurations, as compared with our fiducial configuration. The shaded region indicates the deviation of the values caused by the choice of SPL
Looking beyond only the impact of SPL on derived metallicities, Fig. 12 shows the impact of SED fitting implementations. Given that two of the tested configurations explicitly change the metallicity implementation, it is unsurprising that these have a substantial impact. As expected, the most severe of these comes from assuming a fixed solar metallicity, as most galaxies in the sample would be expected to have sub-solar metallicities. The impact of assuming a constant metallicity means that earlier stars are modelled with a higher metallicity than expected, requiring that the latest stars formed in a galaxy are modelled with a lower metallicity than expected. This is evident in Fig. 12, where it can be seen that the constant Z metallicities are underestimated by around 0.2 dex. Note that the average metallicity across the ZH of a galaxy in our fiducial evolving metallicity model is typically 0.3 dex lower than the fitted Zfinal value, (around Zfinal/2), so the constant Z metallicity is more a reflection of this than the present-day metallicity.
@@ 4.4 Impact on the derived SFHs
Figure 14: The mean Age50 offsets derived when using different SED fitting configurations, as compared with our fiducial configuration. The shaded region indicates the deviation of the values caused by the choice of SPL
@@
While the Age50 provides a simplified way of condensing SFH information, it does not convey the full picture. Fig. 15 shows the median offset in ages computed at different $M_{*,\rm formed}$ percentiles as a function of galaxy stellar mass when using the different SED fitting implementations. As with Fig, 6, is divided into six panels, showing the impact of the individual SPLs.
For a detailed assessment of the behaviour, we first focus on sub-panels (p)–(r), showing the impacts using the ProGeny SPL. In each sub-panel, the black line shows the half-mass age (Age50). Redder lines show the ages computed for lower mass quantiles (corresponding to younger ages, at earlier times) and bluer lines show the ages computer for higher mass quantiles (corresponding to older ages, at later times). Sub-panel (p) shows the relative offset in ages between the constant Z implementation, and the fiducial implementation. For galaxies with $M_{*}<10^{10}\rm{M}_{\odot}$ there is no offset in ages, which explains the global trend in Fig. 14. However, for the most massive galaxies, a deviation is seen. In ProGeny, even though the ages at which galaxies first start forming stars are consistent, the final stars are formed much later in the constant Z implementation than the fiducial implementation, indicating a longer period of star formation. Even though the population-averaged age statistics are consistent, the fact that a bias exists for the most massive galaxies has consequenced in terms of any mass-weighted analysis (such as the inferred cosmic star formation history, see the next Section).
@@
Different behaviour again is noted for the CB19 SPL (sub-panels (m)–(o) of Fig. 15). Like ProGeny the median offsets in the constant Z implementation are biased downwards at high mass, however the colour trend is reversed. Not only are the SFHs skewed towards later times, but the SFH is truncated in terms of duration (such that the early star formation occurs later). In general, the consequences of the $Z_{\odot}$ and b5 SFH implementation are fairly consistent.

### v3

@@ 4.1 Impact on derived stellar masses
The relative deviation in derived stellar masses caused by the choice of SPL is indicated in Fig. 6, where for each SPL we present the mean offset in stellar mass relative to ProGeny. FSPS produces the largest stellar masses, followed closely by ProGeny and CB19, whose masses are only $\sim$0.05 dex lower on average. The other SPLs tend to produce notably lower stellar masses, offset by $\sim$0.1-0.15 dex on average. This behaviour can be directly linked to the underlying stellar recycling fraction implemented in each SPL. As demonstrated by Robotham & Bellstedt (2024), the recycled fraction as a function of stellar age varies between the SPLs. A larger fraction of the remaining stellar mass is accounted for in SPLs with more stellar remnants. FSPS, CB19, and ProGeny all have similar remaining fractions with stellar age, hence explaining why these SPLs tend to recover the most stellar mass. Jones et al. (2022) compared properties of galaxies derived using CB16 (the early form of CB19) and BPASS, and found the CB16 stellar masses to be greater than those from BPASS by 0.14 dex. This is consistent with the offset that we derive between CB19 and BPASS. The open points in Fig. 6 compare the total stellar mass formed throughout the history of the galaxy, which does not account for mass loss due to stellar evolution. It can be seen that the total mass formed by BPASS is more consistent with the other SPLs than the remaining mass, highlighting that the origin of the mass offset comes from the recycled fraction in BPASS.
Figure 7: The median stellar mass offsets derived when using different SED fitting configurations, as compared with our fiducial configuration. The shaded region indicates the mean absolute deviation (MAD) of the masses caused by the choice of SPL (the inter-SPL MAD), where a wider grey bar indicates that the SPL produces more discrepant properties.
@@
The shaded region in Fig. 7 shows the mean absolute deviation (MAD) of the SPL values for the fiducial run, represeting the typical mass deviation created through the SPL selection. What is notable, is that the mass offsets caused by the constant Z and b5 SED configuration changes is less than the offsets simply caused by the choice of SPL.
@@
Figure 8: Stellar mass offsets for individual galaxies using different SED configurations, all compared against the fiducial stellar mass estimates. Top panel shows the impact of assuming a constant metallicity, the middle panel shows the impact of fixing the metallciity to solar, and the bottom panel shows the impact of assuming a stepwise parametric SFH.
While Fig. 7 shows the median stellar mass impact, it is important to note that these discrepancies are stellar mass dependent (as demonstrated in Fig. 8). While the overall offset created by the constant Z assumption is always less than 0.01 dex, at stellar masses $M_{*}>10^{10.5}\rm{M}_{\odot}$ the values are overestimated by 0.06 dex in BPASS, and underestimated by just over 0.04 dex in both CB19 and ProGeny. Subpanels (p)–(r) of Fig. 8 show the offsets of individual galaxies using the ProGeny SPL for each SED configuration, as compared with the fiducial implementation. The impact of ignoring the presence of a relation between a galaxy’s stellar mass and metallicity is evident in sub-panel (q), where the impact of fixing the metallicity to solar is a very strong mass dependence of the stellar mass offset. At low stellar masses, galaxies are forcibly modelled with metallicity values that are too large, forcing the model to produce younger stars to create a sufficiently blue SED. This results in a lower mass-to-light, and therefore the masses are underestimated. At high stellar masses however the assumption of a solar metallicity is more realistic, and hence the stellar mass offset is minimal. Sub-panel (r) of Fig. 8 also shows that while the overall effect of a stepwise versus functional parametrisation is to overestimate the stellar masses by around 0.1 dex in ProGeny, at the highest masses the values are actually consistent. These trends are qualitatively similar for all SPLs except BPASS, where the impact of a constant metallicity is reversed at higher stellar masses (see sub-panel (l) of Fig. 8).
@@ 4.2 Impact on derived SFR
Presenting the results from just the fiducial configuration, we compare the impact of selecting different SPLs on the derived star formation rates (SFRs) relative to the ProGeny-derived SFR in Fig. 9. Within ProSpect, the SFR is taken to be the averaged value of the star formation history in the most recent 100 Myr. Values are generally consistent within ±$\pm 0.1$ dex, where M05 produces the highest values, and BPASS produces the lowest.
Figure 10: The median SFR offsets derived when using different SED fitting configurations, as compared with our fiducial configuration. The shaded region indicates the inter-SPL MAD. All SFR measurements less than $10^{-6}\,\rm{M}_{\odot}\rm{yr}^{-1}$ are shifted to $10^{-6}\,\rm{M}_{\odot}\rm{yr}^{-1}$ for the sake of this comparison (as the SFR measurements derived using massfunc_snorm_trunc can be arbitrarily small).
The relative impact of different ProSpect implementation configurations as compared with the impact of SPL selection is then shown in Fig. 10. As in Fig. 7, the coloured points represent the offset caused by the SED fitting configuration relative to the fiducial configuration, whereas the dark shaded region represents the mean absolute deviation (MAD) of the SPL values for the fiducial run. For runs where the coloured point lies outside of the shaded region, this is an indication that the SED configuration has a greater impact on the SFR measurements than the choice of SPL. The SPL choice is the dominant impact on SFRs in almost all cases. The only exceptions to this are the $Z_{\odot}$ runs conducted with BPASS and M05, where the resulting SFRs are measured to be around 0.2 dex higher than the fiducial implementation, which is on par with the typical impact from SPL choice.
@@
A parameter space that has been extremely well studied in the literature is the stellar mass – SFR plane, which presents the co-called Star Formation Main Sequence (SFMS). Fig. 11 shows how the fiducial implementation produces varying SFMS derivations using the six SPLs. The recovery of the SFMS is generally consistent, with a very similar normalisation, and a similar bending of the MS at high masses. This is reflected by the coloured lines, which represent the running median of all galaxies that have SFRs greater than 1 dex below the SFMS fit by Thorne et al. (2021). While these lines are not meant to be a recovery of the SFMS, they provide a straightforward way of comparing the overall behaviour in the bottom panel. There is a population of high-mass “quenching” galaxies that sit just below the MS in all outputs except that of CB19. While none of the galaxies occupying this region ever appear on the main sequence itself (suggesting that the constraint of this lower SFR is real), there is a tendency for them to scatter to the more fully quenched position below the plot. There is also a population of low-mass “bursting” galaxies produced by M05, however the fact that they are not recovered by any other SPL suggests that the separation to higher SFRs is an artefact of the fitting when using M05. The other SPLs do recover these sources consistently as highly star-forming galaxies, in that they either sit directly on or above the main sequence defined by Thorne et al. (2021). Finally, we note that the derived SFMS all consistently show greater scatter in the b5 SFH implementation than the fiducial implementation, highlighting the consequence of the freedom in the present-day star formation rate when using stepwise bins in the SFH implementation.
@@ 4.3 Impact on the derived metallicity
The derived mass–metallicity relations using different SSPs are given in Fig. 12. Each panel represents the output derived using a different SPL, where values for individual galaxies are indicated as grey points, with the running median shown in the coloured line. The inherent limit in metallicity range provided by each SPL can be seen in this figure, where ProGeny has the widest range in possible metallicity values, whereas M05 has the most truncated range. For a comparison with observations of direct metallicity measurements, we show the Mass – Metallicity Relation (MZR) from Tremonti et al. (2004) as the dashed black line in each panel. The bottom right subpanel compares the median MZR derived by all SPLs. The overall shape for each SPL is similar, recovering the classic increase of metallicity with stellar mass, where there is a break in the relation at $\sim 10^{10}\rm{M}_{\odot}$, and a flattening of the curve at the highest stellar masses.
It is notable that the low-mass values for metallicity are recovered quite differently for the various SPLs. BC03, FSPS, CB19 and ProGeny all have lower median values of metallicity of $\sim 3\times 10^{-3}$. Metallicity values for low-mass galaxies in M05 and BPASS are substantially lower, with $<10^{-3}$. For M05 this is largely driven by the fact that low-mass galaxies are hitting the lower bound of the available templates, while BPASS seems to genuinely prefer fitting lower metallicity values. The SPL that produces a MZR most similar in shape to that of Tremonti et al. (2004) is BC03. It is worth noting that Tremonti et al. (2004) use the BC03 library in the derivation of their metallicity values, so this agreement is perhaps not surprising (while acknowledging that spectrally derived and photometrically derived metallicities are not expected to be the same). The general agreement between photometrically derived and spectrally derived trends using the same SPL and very different techniques is reassuring, however.
@@
Figure 14: The mean Zfinal offsets derived when using different SED fitting configurations, as compared with our fiducial configuration. The shaded region indicates the inter-SPL MAD.
Looking beyond only the impact of SPL on derived metallicities, Fig. 14 shows the impact of SED fitting implementations. Given that two of the tested configurations explicitly change the metallicity implementation, it is unsurprising that these have a substantial impact. As expected, the most severe of these comes from assuming a fixed solar metallicity, as most galaxies in the sample would be expected to have sub-solar metallicities. The impact of assuming a constant metallicity means that earlier stars are modelled with a higher metallicity than expected, requiring that the latest stars formed in a galaxy are modelled with a lower metallicity than expected. This is evident in Fig. 14, where it can be seen that the constant Z metallicities are underestimated by around 0.2 dex. Note that the average metallicity across the metallicity history (ZH) of a galaxy in our fiducial evolving metallicity model is typically 0.3 dex lower than the fitted Zfinal value, (around Zfinal/2), so the constant Z metallicity is more a reflection of this than the present-day metallicity.
@@ 4.4 Impact on the derived SFHs
Figure 16: The mean Age50 offsets derived when using different SED fitting configurations, as compared with our fiducial configuration. The shaded region indicates the inter-SPL MAD.
@@
While the Age50 provides a simplified way of condensing SFH information, it does not convey the full picture. Fig. 17 shows the median offset in ages computed at different $M_{*,\rm formed}$ percentiles as a function of galaxy stellar mass when using the different SED fitting implementations. As with Fig, 8, it is divided into six panels, showing the impact of the individual SPLs.
For a detailed assessment of the behaviour, we first focus on sub-panels (p)–(r), showing the impacts using the ProGeny SPL. In each sub-panel, the black line shows the half-mass age (Age50). Redder lines show the ages computed for lower mass quantiles (corresponding to younger ages, at earlier times) and bluer lines show the ages computed for higher mass quantiles (corresponding to older ages, at later times). Sub-panel (p) shows the relative offset in ages between the constant Z implementation, and the fiducial implementation. For galaxies with $M_{*}<10^{10}\rm{M}_{\odot}$ there is no offset in ages, which explains the global trend in Fig. 16. However, for the most massive galaxies, a deviation is seen. In ProGeny, even though the ages at which galaxies first start forming stars are consistent, the final stars are formed much later in the constant Z implementation than the fiducial implementation, indicating a longer period of star formation. Even though the population-averaged age statistics are consistent, the fact that a bias exists for the most massive galaxies has consequences in terms of any mass-weighted analysis (such as the inferred cosmic star formation history, see the next Section).
@@
Different behaviour again is noted for the CB19 SPL (sub-panels (m)–(o) of Fig. 17). Like ProGeny the median offsets in the constant Z implementation are biased downwards at high mass, however the colour trend is reversed (a feature that also occurs in the M05 SPL). Not only are the SFHs skewed towards later times, but the SFH is truncated in terms of duration (such that the early star formation occurs later). In general, the consequences of the $Z_{\odot}$ and b5 SFH implementation are fairly consistent.
4.4.1 Manifestation of the age–metallicity degeneracy
Figure 18: The distribution of half-mass ages with stellar mass in each SED fitting configuration, with individual SPLs selected. Individual points are coloured by the Zfinal value, and the solid black line in each subpanel shows the running median.
While we have studied the impact of set fitting choices on the metallicity and the ages separately, the known age–metallicity degeneragy means that it is difficult to study these impacts in isolation. Fig. 18 presents the half-mass age as a function of the galaxy stellar mass for each configuration tested, coloured by the final metallicity value Zfinal to better compare these outputs. The overall conclusion from our analysis is that the degree to which the age-metallicity degeneracy falsely drives the fitted parameters changes depends on the SED fitting configuration, which we will elaborate on in this section.
The most immediate impact of the age–metallicity degeneracy noted, is that when the metallicity is set to be fixed in the SED fitting, then the only remaining lever that can be modified to fit the observed colours is age. This can be seen in the $Z_{\odot}$ subpanels, where the resulting age is directly correlated with the galaxy mass (a consequence of the fact that galaxies at higher masses tend to be redder). Interestingly while this behaviour is consistent across the different SPLs, BPASS is a notable exception, as seen by the different appearance of panel (o). The likely cause of this behaviour is the notably different UV flux presence in the BPASS templates. Fig 11 of Robotham & Bellstedt (2024) compiles the templates at solar metallicity across a range of ages for each of the SPLs studied in this work. At every age, the BPASS template is an outlier within the UV. At lower ages the UV flux is higher, and at older ages the UV flux is lower. This may well explain the slightly bimodal behaviour in panel (o) of Fig. 18, where at fixed $Z_{\odot}$ ProSpect tends to choose either younger or older ages, depending on the UV flux in a galaxy.
When metallicity is allowed to vary, then at a given stellar mass the range in derived ages becomes greater. This represents a relaxing of the consequence of the age–metallicity degeneracy. In the constant Z implementation, the sample-wide range in recovered ages tends to be lower than in the $Z_{\odot}$ implementation, showing that the freedom in metallicity is a more dominant lever when trying to reproduce the observed colours of a galaxy. The fiducial implementation, with its evolving metallicity prescription, presents a compromise between these two trends. The variation of age across each mass bin is still broad, highlighting that the age lever is not dominating, however the sample-wide range in recovered ages is also broader, highlighting that the fitted metallicity is facilitating the more extreme age values that would be hampered by a constant metallicity implementation. At the highest masses, the recovered ages are older in the fiducial panel than the constant Z panel, and have a higher Zfinal value. This is because the oldest stars can form at a low metallicity in the fiducial case, allowing ProSpect to favour the oldest ages. Again, a notable exception to this trend is BPASS, where the highest mass galaxies in the constant Z implementation in panel (n) actually have older ages than the same galaxies in the fiducial implementation in panel (m). We believe this can again be attributed to the UV flux behaviour in BPASS, where for the highest metallicities, BPASS consistently produces more UV flux than other SPLs (see fig B18 in Robotham & Bellstedt 2024). This leads ProSpect to disfavour high-metallicity outputs in the fiducial implementation, thereby limiting the ages somewhat.
This analysis highlights that when limited flexibility is afforded in an SED fitting implementation, the age–metallicity degeneracy will drive the resulting properties. When afforded greater flexibility and more physical implementations however, the impact of this degeneracy is minimised.
Finally, evident from the b5 subpanels in Fig. 18 is the discretisation in age, reflecting favoured half-mass age values when using the stepwise parametric function for the SFH. This will occur when a large fraction of the stellar mass in a galaxy is formed within a single time bin, which because of the fixed width of each bin results in discretised half-mass ages. The existence of these artefacts in the age derivation mean that the impact of the age-metallicity degeneracy here is more difficult to discern.